sugermomma:

Together, teammates Pedro and Ricky 2677 got base hits last season. Pedro 281 had more hits than Ricky. How many hits did each player​ have?

sugermomma:

what will it give me

Krissy525:

welcome to qc

LostLegacy06:

You have to turn this into an expression

LostLegacy06:

x + 281=2677, should help you find your answer

Krissy525:

let x equal the amount of hits Ricky has x+281- pedro x- ricky x+281+x=2677 2x+281=2677 subtract 281 from both sides 2x=2396 divide by 2 x=1198 ricky has 1198 pedro has 1479